Hello,
I have a question I need help with. Your time is appreciated.
From a point on a straight road, two cars are driven in opposite directions, one at 47 mph and the other at 51mph. In houw many hours will they be 392 miles apart?
Thank you

Found 2 solutions by rfer, sudhanshu_kmr:
Answer by rfer(16322)

moving apart at 98 mph
392/98=4 hrs

Answer by sudhanshu_kmr(1152)


In one hour they will apart by ( 47 + 51 )= 98 miles
time = 392/98 = 4
so, 4 hours.